OpenStudy (anonymous):

Hi everyone! Can someone check my work for this infinite series solution to the DE y"=xy I am supposed to find y1 and y2. Any advice would be welcome. Thanks! :o)

OpenStudy (anonymous):

At first I was concerned about not having the values of Co and C1, but somehow just following y=Co+C1x+C2x^2 etc, it seemed to work itself out...If anyone cares to elaborate on how it did that , that would be great! :o)

OpenStudy (amistre64):

so your trying power series solution eh

OpenStudy (anonymous):

yes...ugh! :o)

OpenStudy (amistre64):

y = sum0 an x^n xy = sum0 an x^{n+1} y'' = sum2 an n(n-1) x^{n-2} line up the powers, i work it to the lowest, so n-2 xy = sum0+3 a[n-3] x^{n-3+1} y'' = sum2 an n(n-1) x^{n-2} xy = sum3 a[n-3] x^{n-2} y'' = sum2 an n(n-1) x^{n-2} line up the indexes xy = sum3 a[n-3] x^{n-2} y'' = 2a[2] + sum3 an n(n-1) x^{n-2} combine the an coeefs since it all lines up y'' - xy = 2a[2] + sum3 {an n(n-1) - a[n-3]} x^{n-2} = 0 we good to here?

OpenStudy (anonymous):

did you look at my work?

OpenStudy (amistre64):

a0 = a0 a1 = a1 a2 = 0 i did, but i tend to have to work this out to make sure it flows ... lets see if we end up the same

OpenStudy (anonymous):

it's much easier to refer to what I already did instead of trying to figure out all the typing and syntax

OpenStudy (amistre64):

im up to you on the bottom of the first page

OpenStudy (anonymous):

gotcha! :o) just so you know though, it's a little hard for me to follow the typing fyi

OpenStudy (amistre64):

our rule for the coeffs an n(n-1) - a[n-3] = 0 a[n] = a[n-3]/(n(n-1)) \[a_n=\frac{a_{n-3}}{n(n-1)}~:~n\ge3\] we have the same setup so far

OpenStudy (amistre64):

why do we need 4 non-negative terms?

OpenStudy (anonymous):

because of the instructions

OpenStudy (amistre64):

and it tends to be best if you can adjust your denominators to be factorials ... or at least thats the way i work them

OpenStudy (anonymous):

i think I accidentally did 5...opps

OpenStudy (anonymous):

well some of the denominators are not factorials though...some are like: 9.8.6.5.3.2 they are not 9.8.7.6.5.4.3.2.1 etc they skip so I can't write as factorials unless you know something that I don't

OpenStudy (amistre64):

you can make them factorials by multiply by some useful form of 1 to fill in what they need \[a_n=\frac{a_{n-3}}{n(n-1)}~:~n\ge3\] \[a_3=\frac{a_{0}}{3(2)}=a_0\frac{1}{3!}\] \[a_4=\frac{a_{1}}{4(3)}=a_1\frac{2}{4!}\] \[a_6=\frac{a_{3}}{6(5)}=a_0\frac{1(4)}{6!}\] \[a_7=\frac{a_{4}}{7(6)}=a_1\frac{2(5)}{7!}\] \[a_9=\frac{a_{6}}{9(8)}=a_0\frac{1(4)(7)}{9!}\] \[a_10=\frac{a_{7}}{10(9)}=a_1\frac{2(5)(8)}{10!}\]

OpenStudy (amistre64):

our arbitrary constants are ...well you have c's. c0 and c1

OpenStudy (amistre64):

\[y_1=a_{3n}=a_0\frac{\prod\limits_{k=1}^{n}(3k-2)}{(3n)!}\] \[y_1=a_{3n+1}=a_1\frac{\prod\limits_{k=1}^{n}(3k-1)}{(3n+1)!}\]

OpenStudy (amistre64):

you as long as all your computations on the coeffs are right, it looks like you did a fine job

OpenStudy (anonymous):

my teacher doesn't want the answers in series form, just y1 and y2 expanded

OpenStudy (amistre64):

you did fine ...

OpenStudy (anonymous):

okay great! :o) do you mind if I ask a couple things about your method?

OpenStudy (amistre64):

sure

OpenStudy (amistre64):

i mean go ahead and ask lol ...

OpenStudy (anonymous):

you typed in that n>=3 thingy my teacher did that too but I didn't know why I thought I heard him say he just switched back the index to n= instead of k= because he said it's easier to calculate do you think I heard him correctly? I mean, is that why you changed your index back as well?

OpenStudy (amistre64):

ive tried it both ways, and im more comfortable (learned it that way) defining a_n= instead of a_(n+3)= when we drop the exponents down to the 'lowest' value, the n-a to me reads as the lowest value of a is the largest number. that allows me to move my index up instead of down. its a habit. now because when i lined all the powers and indexes up, my summation started at 3, thats why i defined n to start at 3 a_(n-3) after all only makes sense if we start with n=3

OpenStudy (anonymous):

ok, so my teacher(or you for that matter) totally didn't need to switch back the indices from k= back to n= then...it's simply a function of your comfort level and preference then right?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

okay, well I better stick with k= I think, the less switching for me the better...algebra hates me! :o~

OpenStudy (amistre64):

:) potato tomato eh

OpenStudy (anonymous):

lol...okay something else...

OpenStudy (anonymous):

I completely understand why you would want to represent the denominators with factorials so that you can represent the y1 and y2 as power series representations... so... taking my C9 on page 2 for example, how on Earth do I do that? C9=Co/9.8.6.5.3.2

OpenStudy (amistre64):

you are missing 1 and 4 and 7 underneath (the 1 is useless i know but habit) multiply by 4(7)/4(7)

OpenStudy (anonymous):

gimme a sec to write that out

OpenStudy (anonymous):

so 28Co/9! ?

OpenStudy (amistre64):

Co 28/9! yes

OpenStudy (anonymous):

yikes! looks scary

OpenStudy (amistre64):

spose we had something like: 1/(7.3.2) we are missing 6.5.4 1/(7.3.2) = (6.5.4)/(7.6.5.4.3.2) = 120/7!

OpenStudy (anonymous):

yeah I think I see that now... I was just noticing that my next missing term in the denominator would have been C10, so if I needed 3 terms in the denominator to make a factorial, I could have just done: Co/9.8.6.5.3.2 times (4/7)(7/4)(1/10)(10/1) = 280Co/10! OR... Co/9.8.6.5.3.2 times (4/7)(7/4)(10/10)= 280Co/10! Is that correct? :o/

OpenStudy (amistre64):

the thing is, by filling in the factorials, we notice a pattern of products forming up top for a given nth term 1 1(1+3) 1(1+3)(1+3+3) 1(1+3)(1+3+3)(1+3+3+3) product (k=1 to n) of: 1 + 3(k-1) or product (k=1 to n) of: 3k - 2 for some nth term

OpenStudy (amistre64):

\[c_{10}=c_1\frac{1}{10.9.7.6.4.3}=c_1\frac{\color{red}{8}.\color{red}{5}.\color{red}{2}}{10.9.\color{red}{8}.7.6.\color{red}{5}.4.3.\color{red}{2}}\] \[~~~~~~~~~~~~\implies c_1\frac{80}{10!}\]

OpenStudy (anonymous):

yes, I should have known it wasn't that simple, I never did the math to figure out C10 but regardless, I get it! A more simple approach would be this: "Hey there Ms. math student, are you missing 3 6 and 8 from the denominator? No problem, just multiply the quantity by (3/3)(6/6)(8/8)" Right? :o)

OpenStudy (amistre64):

correct :)

OpenStudy (anonymous):

lol...last thing if you don't mind...

OpenStudy (anonymous):

at the beginning you stated right away that Co=Co and C1=C1 now for me I was a bit worried just from lack of experience that I didn't have values for Co and C1, but am I correct in assuming that as long as I put everything in terms of Co and C1 and then factor out at the end Co and C1, I never never actually need to state that Co=Co and C1=C1 do I?

OpenStudy (amistre64):

before you learnt power series, you dealt with arbitrary constants .... the C0 and C1 are nothing more than the arbitrary constants. If we had an initial value we might even be able to solve for them.

OpenStudy (amistre64):

the wolf gives us the solution as \[y=c_0Ai(x)+c_1Bi(x)\] Ai is the Airy function Bi is the Airy Bi function i got no idea what those are, but they are what they call the power series that we generated.

OpenStudy (anonymous):

so then I don't need to worry about them since I am guided by: y=Co+C1x+C2x^2 etc and I put all the C2 AND c3'S etc in terms of C2 and C3 etc they get factored out, leaving your y1 and y2, and hence no need to be a scared-E pants right? :o)

OpenStudy (anonymous):

sorry...typo

OpenStudy (amistre64):

no need to be concerned :)

OpenStudy (anonymous):

so then I don't need to worry about them since I am guided by: y=Co+C1x+C2x^2 etc and I put all the C2 AND c3'S etc in terms of Co and C1 etc they get factored out, leaving your y1 and y2, and hence no need to be a scared-E pants right? :o)

OpenStudy (anonymous):

I think I got it finally! Any further advice for me?

OpenStudy (amistre64):

dont sweat the small stuff ....

OpenStudy (anonymous):

:o) Thanks sooooo much @amistre64 ! You definitely get a medal! They should have a cookie button on here as well! nomnomnom :o) Thanks!

OpenStudy (amistre64):

>(B)< ; butterscotches lol

OpenStudy (anonymous):

omg...butterscotch topping on french vanilla ice cream! mmm

OpenStudy (anonymous):

I think I need a break now! whew no ice cream but I do have M and M's meh! lol Have a great night!

OpenStudy (amistre64):

imma leave before one of us goes into a diabetic coma ... good luck ;)

OpenStudy (anonymous):

Thanks you too